[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Sep 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Sep 29 03:17:05 UTC 2017


HSU Shift Log for Thu Sep 28 2017 23:16:58 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: NCEP model guidance (GFSI, HWFI, HMNI) were not available at  
00z. Monica fixed a script that was modified today and they models were  
available before the 03z advisories had to be finished.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170929/9d1d4feb/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list